Jen bought 4 tickets.Amber bought 5 tickets.The tickets cost $2 each.How much did the girls spend on tickets in all?
1)What is the hidden question you need to answer before you can solve the problem
2)Solve the problem. Complete the bar diagrams.Show the equations you used
3)How would your equations change if Amber bought only 3 tickets? Explain

Found 2 solutions by ikleyn, richwmiller:
Answer by ikleyn(53763)

.

The problem is solved in ONE LINE


    The girls spent  4*2 + 5*2 = 8 + 10 = 18 dollars, in all.


All the rest in the post does not deserve any attention.


I do not like VERY MUCH the form on how this text and instructions in the post are organized.

They only suppress the young students' mind.

In my view, it is very bad style teaching young students.


/\/\/\/\/\/\/\/

By the way, the problem is about arithmetic, not about Algebra.

So, it is unclear to me, why you posted it to this forum.



Answer by richwmiller(17219)

Although this website is named algebra.com, any problems that involve any type of mathematics, whether it be simple arithmetic or advanced algebra, is welcome on this website.
Any physics, chemistry, biology, finance or statistics problems that involve any type of math are welcome here.